state the equation( in slope-intercept form) of the line that passes through the points (-2,5) and (4,8)
Answer by Alan3354(69443)

state the equation( in slope-intercept form) of the line that passes through the points (-2,5) and (4,8)
------------
Find the slope, m
m = diffy/diffx = (5-8)/(-2-4) = -3/-6
m = 1/2
---------
Then use y=mx+b with either point
(4,8)
8 = (1/2)*4 + b = 2+b
b = 6
--> y = (1/2)x + 6
------------
------------
Check both points:
5 =? (1/2)(-2) + 6
5 = 5
------------
8 =? (1/2)*4 + 6
8 = 8